In the past, a company used to produce 40 tons of a certain food per day on average. Now, using a new technology, the mean da1b)The value of the standardized test statistic is:

Group of answer choices

a) 1.00

b) 5.00

c) None of the above

d) -5.00

e) -1.00

1c)Find the rejection region and state your conclusion at \alphaα = 0.05.

Group of answer choices

a) Reject region: t < -1.711; Decision: Fail to reject H0

b) Reject region: z > 1.645; Decision: Reject H0

c) Reject region: t < -2.064 or z > 2.064; Decision: Reject H0

d) Reject region: t > 1.711; Decision: Reject H0

e) Reject region: z < -1.96 or z > 1.96; Decision: Fail to reject H0
Answer #1

a) Option-C) H0 : µ < 40 versus H1 : µ > 40

b) Test statistic =\frac{\bar{X}-\mu}{s/\sqrt{n}}

                        =\frac{45-40}{5/\sqrt{25}}

                        = 5

Option-B) 5

c) t0.05,24 = 1.711

Option-D) Reject region: t > 1.711; Decision: Reject H0